OpenStudy (lukecrayonz):

log(4x-3)+6=4 log(4x-3)+6=4 @Mathematics

OpenStudy (mertsj):

10^-2 = 4x-3

OpenStudy (anonymous):

log(4x-3) = -2 10^-2=4x-3 1/100 = 4x-3 301/100=4x x=301/400